2tan’3/4 = ? ans sin’24/25 prove me i m not able to solve this

2tan’3/4 = ?
 
ans sin’24/25 
prove me i m not able to solve this

Grade:12

1 Answers

ritesh
7 years ago
see...2tan-1 x=sin-1{2x/1 + x2} ...- (i)
 
given 2tan-1(3/4)...here x = 3/4
 
on substituting value of  x in (i) we get...
RHS
sin-1 {2 * (3/4) / 1 + (3/4)2} ….which on solving gives 
2tan-1¾ = sin-1(24/25)
hope u got the method of solving this